The school board estimates that there will be a 2.5% increase in enrollment for next year. If the present enrollment is 8,657, what is the estimated enrollment?
Found 2 solutions by wgunther, jayanthihemkos@yahoo.com:
Answer by wgunther(43)

2.5% is .025, so a 2.5% increase in something means multiply that something by (1+.025).

Answer by jayanthihemkos@yahoo.com(87)

The school board estimates that there will be a 2.5% increase in enrollment for next year. If the present enrollment is 8,657, what is the estimated enrollment
Present enrollment = 8,657
Enrollment next year = 8,657+2.5% of 8,657
= 8,657+2.5/100*8,657
= 8,657+216.425
= 8873.425
The estimated enrollment = 8873